OpenStudy (jacobbirddog):

Anyone good with finding area? if so i would like some help... fan+medallion :)

OpenStudy (anonymous):

let's see if i can help, type the problem

OpenStudy (jacobbirddog):

im stumped :P

OpenStudy (jacobbirddog):

theres the problem

OpenStudy (anonymous):

Divide the shape into three parts.  Area 1 has the following area. \[A_1 = 14 \cdot 16\] Area 2 has the following area. \[A_2 = 54 \cdot (28 - 16)\] Area 3 has the following area. \[A_3 = {1 \over 2} (54 - 22 - 14) \cdot (28 - 16)\] Then you just add them up to get the total area.

OpenStudy (anonymous):

28*14+22*(28-16)+(54-22-14)*(28-16)+16/2*(54-22-14) = 1016

OpenStudy (jacobbirddog):

fanned both of ya, medallion goes to the texas tech due to an explination :)

OpenStudy (jacobbirddog):

sorry for bad english, it's not my first language

OpenStudy (anonymous):

the approach is always the same decompose the figure into squares and triangles of know dimension and sum their areas another fast approach is this one: 28*54-(54-14+22)*8 bigger square - trapezoid: